# Evaluating and Improving the Effectiveness of Synthetic Chest X-Rays for Medical Image Analysis

### v1

@@ Results
However, we found that limited radiologist feedback is minimally effective in refining model outputs due to the limited size of the expert cohort providing preferences.
Despite these advancements, the study has some limitations. Our evaluation focused exclusively on chest X-rays, which limits understanding of how well the approach works for other imaging modalities. The tasks we explored were also narrow, as segmentation experiments only targeted pneumothorax. Additionally, relying on a small group of radiologists to collect preference data for fine-tuning may have constrained the model’s optimization, reducing its generalizability. Expanding this work to include diverse imaging types, a broader range of tasks, and feedback from more radiologists could strengthen the effectiveness and applicability of this approach.

### v2

@@ 5 Conclusion
However, we found that expert feedback is minimally effective in refining model outputs due to the limited size of the expert cohort providing preferences.
Despite these advancements, the study has some limitations. Our evaluation focused exclusively on chest X-rays, which limits understanding of how well the approach would work for other imaging modalities. The tasks we explored were also narrow, as segmentation experiments only targeted pneumothorax. Additionally, relying on a small group of radiologists to collect preference data for fine-tuning may have constrained the model’s optimization, reducing its generalizability. Expanding this work to include diverse imaging types, a broader range of tasks, and feedback from more radiologists could strengthen the effectiveness and applicability of this approach.